OpenStudy (anonymous):

Hey! can I do anything with -5 = -2^x. Find x? Thanks :)

OpenStudy (anonymous):

or 2^x = 5?

OpenStudy (doc.brown):

I don't know, maybe\[2^{m+n}=5\]\[2^m\times2^n=5\]\[2^2+2^n=5\]Somehow get \(2^n\) to equal 1.25

OpenStudy (anonymous):

I changed the second one (2^x =5) to logs (ie. log2(5) = x). and got approx. 2.3219. But I don't know whether we can use our calculators or not so I was just wondering if there was another way.